A tugboat can pull a boat 24 mi downstream in 2h.going upstream the tugboat can pull the same boat 16 mi in 2h. what is the speed of the tugboat in still water ? what is the speed of the current ?
how can use the formula d=rt to help you solve the problem?
how are the tugboat's speed when traveling upstream and downstream related to its speed in still water and the spend of the current?

Answer by mananth(16946)

Boat speed =x mph
current speed =y mph
against wind x-y mph 2 hours
with wind x+y mph 2 hours

Distance upstream= 16 miles Downstream = 24 miles
t=d/r against current
16/(x-y )=2
2(x-y) =16
2x-2y = 16 ....................1

24/(x+y )= 2.00
2.00(x +y)= 24.00
2.00x+ 2.00 y =24.00 ...............2
Multiply (1) by 1
Multiply (2) by 1
we get
2 x + 2 y = 16
2 x + 2 y = 24
4 x = 40
/ 4
x = 10.00 mph

plug value of x in (1)
2.00 x -2.00 y = 16.00
20.00 -2.00 y = 16.00
-2.00 y = 16.00 -20.00
-2.00 y = -4.00
y = 2 mph
Boat 10 mph
current 2 mph